At 1000 K, a sample of pure NO2 gas decomposes:

2NO2(g) ⇌ 2NO(g) + O2(g)

The equilibrium constant, KP, is 158. Analysis shows that the partial pressure of O2 is 0.11 atm at equilibrium. Calculate the pressure of NO and NO2 in the mixture.

Pressure of NO:

__atm

Pressure of NO2:

__atm
Answer #1

-------------- 2NO2(g)     ⇌    2NO(g) +    O2(g)

I--------------- P ----------------   0 -------------   0

C-------------- -2x --------------- +2x ------------ +x

E------------ P-2x -------------- 2x -------------- x

at equilibrium partial pressure of O2   PO2   = x = 0.11atm

at equilibrium partial pressure of O2   PNO   = 2x = 2*0.11 = 0.22atm

Kp   = P^2NO*PO2/P^2NO2

158    = (0.22)^2*0.11/P^2NO2

P^2NO2   = (0.22)^2*0.11/158

P2NO2   = 3.37*10^-5

PNO2    = 0.0058atm